OpenStudy (anonymous):

Is there anybody on the face of this planet who is good at calculus and can help me until I am done with all my study sheets?!?!?!

OpenStudy (anonymous):

I got sick and have finals and I need to get all of my study sheets done so I am prepared for my finals.

OpenStudy (isaiah.feynman):

I can't spend all day here, because of my own exams. But I could help for a bit and hopefully other will continue from where I stopped.

OpenStudy (anonymous):

use limits to determine if \[f(x)=\ln (x-1)\] is continuous at x=1.

OpenStudy (anonymous):

i do have options. \[A. \lim_{x \rightarrow 1+}f(x)=\lim_{x \rightarrow x-}f(x)=f(1)\] so the function is continuous at x=1 \[B. \lim_{x \rightarrow 1+}f(x)\neq \lim_{x \rightarrow 1-}f(x)\]so the function is not continuous at x=1

OpenStudy (anonymous):

\[C. \lim_{x \rightarrow 1+}f(x)=\lim_{x \rightarrow 1-}f(x)\neq f(1)\] so the function is not continuous at x=1 \[D.\]can not determine using limits

OpenStudy (phi):

when x ->1+ it means x is just a bit bigger than 1, say 0.001 as a concrete example what is ln( 1.001 - 1) ? now compare to x-> 1-, which means x is a bit less than 1: what is ln( 0.999 - 1) ?

OpenStudy (anonymous):

-6.90775528 + 3.14159265 I???

OpenStudy (phi):

you can use a calculator. You will see you get -6.9 for the first case and "math error" for the second. That means they are not the same

OpenStudy (phi):

or you get a complex number. but definitely different.

OpenStudy (anonymous):

so it would be B?

OpenStudy (phi):

I think so.

OpenStudy (anonymous):

An object in free fall has its distance from the ground measured by the function d(t)=-4.9t^2 +50, where d is in meters and t is in seconds. If gravity is the only acceleration affecting the object, what is gravity's constant value (include units in your answer)?

OpenStudy (phi):

if you use x instead of d (to avoid confusion) \[ x= -4.9t^2 +50\] by definition \[ \text{velocity }= \frac{dx}{dt} \\ \text{acceleration }= \frac{d^2x}{dt^2} \]

OpenStudy (anonymous):

but what in the equation is d and what is x and t?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

d is just an operator to represent derivatives (rates of change) x is the horizontal distance t is time

OpenStudy (anonymous):

so it is the first derivative of x divided by the first derivative of time?

terenzreignz (terenzreignz):

no, dx/dt is THE derivative of x with respect to time.

OpenStudy (isaiah.feynman):

Sorry to interrupt guys, @kaylalynn have you learned differentiation?

OpenStudy (anonymous):

I was sick for about 3 weeks ND my teacher just gave me these to get ready for finals. im so lost. so here I am.

OpenStudy (isaiah.feynman):

I am so sorry. I think your teacher can do better.

OpenStudy (anonymous):

I agree. however this is all I got to even prepare me . :(

OpenStudy (anonymous):

Do you know the definition of differential?

OpenStudy (anonymous):

definition? no

OpenStudy (isaiah.feynman):

No, she missed many classes how can she know. I think your teacher should give a little more time.

OpenStudy (anonymous):

I don't blame her completely. she needs to have final scores in by Tuesday.

OpenStudy (isaiah.feynman):

The challenge now is how folks here will TEACH you 3 weeks work in a couple of hours. Cos we can't just solve and explain

OpenStudy (anonymous):

yea I know cuz then I wouldn't be ready for finals:(

OpenStudy (anonymous):

all I can think of is if somebody could answer and explain and ill try to keep up. If I get lost, I can say something??

OpenStudy (phi):

When (or if) you have time, you can learn everything watching Khan's videos This is an intro type video. http://www.khanacademy.org/math/calculus/differential-calculus/derivative_intro/v/calculus--derivatives-1--new-hd-version

OpenStudy (anonymous):

okay thank you. I can try that.

OpenStudy (phi):

For your question *** An object in free fall has its distance from the ground measured by the function d(t)=-4.9t^2 +50, where d is in meters and t is in seconds. *** I renamed the distance d(t) as x(t) because calculus uses d to stand for differential if x is distance (that changes as time time changes) the change in distance divided by change in time is speed (or velocity) we say \[ V= \frac{dx}{dt} \] where \[ x(t) = -4.9t^2 +50 \] take the derivative of both sides \[ \frac{dx}{dt} = \frac{d}{dt}(-4.9t^2 + 50 ) \]

OpenStudy (phi):

to do the derivative , you need to know the short-cut rules Here is one http://www.khanacademy.org/math/calculus/differential-calculus/power_rule_tutorial/v/power-rule

OpenStudy (anonymous):

is it 9.8m/s^2????

OpenStudy (phi):

yes

OpenStudy (phi):

you should find \[ \frac{dx}{dt} = \frac{d}{dt}(-4.9t^2 + 50 ) \\ = -4.9\frac{d}{dt}t^2 + \frac{d}{dt}50 \\ -4.9 \cdot 2 t + 0\\ = -9.8 t \] now take the derivative again \[ \frac{d}{dt} \left(\frac{dx}{dt} = -9.8t\right) \\ \frac{d^2x}{dt^2} =-9.8 \frac{dx}{dt} t\\ \frac{d^2x}{dt^2} =-9.8 \frac{m}{s^2} \]

OpenStudy (phi):

Here are youtube links. Do they work? Intro https://www.youtube.com/watch?v=ANyVpMS3HL4 Power rule https://www.youtube.com/watch?v=bRZmfc1YFsQ

OpenStudy (anonymous):

how do I find the equation of a tangent line using a table?

OpenStudy (isaiah.feynman):

Oops I've not heard of that before.

OpenStudy (phi):

can you post the exact question ?

OpenStudy (anonymous):

The table below gives specific values to a differentiable function h(t). Use this information to write an equation for the tangent line to the graph h(t) at t=9.

OpenStudy (anonymous):

t 0 5 9 14 20 f(t) 86 77 64 59 52

OpenStudy (phi):

my only thought is to find the slope through the points (5,77) and (14,59) and write the equation of a line through the point (9,64)

OpenStudy (anonymous):

slope is -2?

OpenStudy (phi):

yes, now use y - 64 = -2(x-9)

OpenStudy (phi):

or y= -2x +82

OpenStudy (phi):

But I would like to see your notes/write-up on how to do this... I am only guessing on this problem

OpenStudy (anonymous):

I got y=-2x+192?

OpenStudy (phi):

y= -2x + 18 + 64 y= -2x + 82

OpenStudy (anonymous):

\[y-64=-2(x-64)\] \[y-64=-2x+128\] \[y-64+64=-2x+128+64\] \[y=-2x+192\]

OpenStudy (phi):

Oh, you are a victim of my typo we should be using the point (9,64) and I accidentally used 64 instead of 9 the equation of a line through the point (x0,y0) is y - y0 = m(x - x0) Here x0 is 9 and y0 is 64

OpenStudy (phi):

Do you have *any* notes on how they would do this problem ?

OpenStudy (anonymous):

no I don't and that equation doesn't work for the rest...

OpenStudy (phi):

No the equation will only works (if it works at all) for one point on the curve. Your points, if you plot them do not line up... we have a curve that changes direction. A tangent line at (9,64) is expected to touch the curve only at (9,64) A tangent line at another spot on the curve will have a different slope (if the curve changes direction... and it does)

OpenStudy (anonymous):

I have a funeral viewing I have to go to . ill be back later tonight

OpenStudy (anonymous):

did we figure anything out?